# Dominated Actions in Imperfect-Information Games

### v6

@@ 7 Conclusion
Dominance is a fundamental concept in game theory. It is well-understood in normal-form games, but its impact in imperfect-information games has so far been unexplored. We consider several plausible definitions of dominated actions which we demonstrate to be problematic; however, one of them which we denote as strong domination can be useful as an efficient preprocessing step. We present a new definition that addresses these limitations. We show that both strictly and weakly dominated actions can be identified in polynomial time, and that iterative removal can be performed in polynomial time in $n$-player games. Our algorithms identify actions that are dominated by any mixed strategy, not necessarily a pure action. We demonstrate that removing dominated actions can play a significant role in reducing the size of realistic imperfect-information games. This can serve as an efficient preprocessing step before computation of a Nash equilibrium. Subsequent recent work has shown that removal of dominated actions enabled an algorithm to compute a Nash equilibrium in a three-player imperfect-information game in under 3 seconds while the algorithm was unable to solve the full game in 24 hours [3]. In practice our algorithms could be sped up by several heuristics such as selection of the order of traversal of information sets and players. Recent work has shown that some games contain many “mistake” actions that are played with probability zero in all Nash equilibria but are not dominated [2]. Thus, there is potentially more future ground to explore on efficient game reduction by elimination of poor actions.

### v7

@@ 6 Conclusion
Dominance is a fundamental concept in game theory. It is well-understood in normal-form games, but its impact in imperfect-information games has so far been unexplored. We consider several plausible definitions of dominated actions which we demonstrate to be problematic; however, one of them which we denote as strong domination can be useful as an efficient preprocessing step. We present a new definition that addresses these limitations. In games with perfect recall and publicly observable actions we show that both strictly and weakly dominated actions can be identified in polynomial time, and that iterative removal can be performed in polynomial time in two-player games. Our algorithms identify actions that are dominated by any behavioral strategy, not necessarily a pure action. We demonstrate that removing dominated actions can play a significant role in reducing the size of realistic imperfect-information games. This can serve as an efficient preprocessing step before computation of a Nash equilibrium. Subsequent recent work has shown that removal of dominated actions enabled an algorithm to compute a Nash equilibrium in a three-player imperfect-information game in under 3 seconds while the algorithm was unable to solve the full game in 24 hours [3]. In practice our algorithms could be sped up by several heuristics such as selection of the order of traversal of information sets and players. Recent work has shown that some games contain many “mistake” actions that are played with probability zero in all Nash equilibria but are not dominated [2]. Thus, there is potentially more future ground to explore on efficient game reduction by elimination of poor actions. The most immediate open problems for future study are to determine the complexity of finding dominated actions in games in which the assumptions of publicly observable actions and/or perfect recall do not hold, as well as studying games with $n>2$ players.